# biblatexref.bib

### v2

@@ 1. Introduction
is satisfied, where $g(C)$ is the geometric genus of $C$. The goal of this paper is to prove the algebraic hyperbolicity of very general hypersurfaces in homogeneous varieties by generalizing techniques developed by Clemens [Cl86, Cl03], Ein [Ei88, Ei91], Voisin [Vo96, Vo98], Pacienza [Pa03, Pa04], Clemens and Ran [ClR04], Coskun and Riedl [CR19a, CR19] and Yeong [Y22].
A complex manifold $X$ is called (Brody) hyperbolic if every entire map $f:\mathbb{C}\to X$ is constant, and it is Kobayashi hyperbolic if the Kobayashi pseudometric is nondegenerate. In general, Kobayashi hyperbolicity implies Brody hyperbolicity, and when $X$ is a compact complex manifold, Brody proved that the converse holds [Br78]. Hyperbolicity is conjectured to have deep connections to the geometry and arithmetic of varieties:
@@
See Demailly’s survey paper [De20] for an exposition on hyperbolicity.
It can be, however, challenging to prove the hyperbolicity of a given variety. Demailly introduced algebraic hyperbolicity as an algebraic analogue for hyperbolicity, and proved that, for smooth projective varieties, hyperbolicity implies algebraic hyperbolicity and conjectured the converse [De95].
A lot of progress has been made in the study of the algebraic hyperbolicity of very general hypersurfaces in $\mathbb{P}^{n}$. For $n=3$, Xu [Xu94] proved that very general hypersurfaces $X\subset\mathbb{P}^{3}$ of degree $d\geq 6$ are algebraically hyperbolic. This was improved by Coskun and Riedl [CR19a], who showed that also a very general quintic surface is algebraically hyperbolic. Since surfaces of degree at most four contain rational curves and cannot be algebraically hyperbolic, the classification for very general hypersurfaces in $\mathbb{P}^{3}$ is complete. For $n\geq 4$, Clemens and Ein [Cl86, Ei88] proved that $X$ is algebraically hyperbolic when $d\geq 2n$. This was improved to $d\geq 2n-1$ by Voisin [Vo96, Vo98]. Pacienza [Pa04] and Clemens and Ran [ClR04] proved it for $d\geq 2n-2$ and $n\geq 6$. Yeong [Y22] did the octic 4-fold case and improved the result to $d\geq 2n-2$ and $n\geq 5$. If $d\leq 2n-3$, $X$ contains lines, thus it is not algebraically hyperbolic. Hence, the only remaining case in $\mathbb{P}^{n}$ is for sextic threefolds.
Haase and Ilten [HI21] started to classify algebraically hyperbolic surfaces in toric threefolds, leaving a list of open cases at the end of their paper. Coskun and Riedl resolved almost all of their open cases in [CR19]. Robins [Rob23] uses Haase and Ilten’s method to determine many algebraically hyperbolic surfaces in toric varieties with Picard rank 2 and 3. Yeong [Y22] applies Coskun and Riedl’s techniques to almost completely classify algebraically hyperbolic hypersurfaces in $\mathbb{P}^{m}\times\mathbb{P}^{n}$, except for some degrees in $\mathbb{P}^{3}\times\mathbb{P}^{1}$.
@@
Let $A$ be a smooth complex variety such that:
  • $A$ is homogeneous with a transitive action by an algebraic group $G$,
Write the class of the canonical divisor of $A$ as $K_{A}=a_{1}H_{1}+\cdots+a_{m}H_{m}$. Let $\mathcal{E}$ be the line bundle $d_{1}H_{1}+\cdots+d_{m}H_{m}$ for $d_{1},\ldots,d_{m}>0$, invariant under $G$, and assume that $H_{1},\ldots,H_{m}$ is a collection of section-dominating line bundles for $\mathcal{E}$ (see section 2). A degree $(d_{1},\ldots,d_{m})$ hypersurface $X$ of $A$ is a section of $\mathcal{E}$.
@@
The case $d_{i}\leq\dim A-a_{i}-3$ for some $1\leq i\leq m$ is left open. When $\dim A\leq 4$, there are examples when it fails to imply algebraic hyperbolicity (see Examples 4.4 and 4.5.) For $A=\mathbb{P}^{4}$, the case of sextic threefolds remains open. However, when $\dim A\geq 5$, it does imply algebraic hyperbolicity for $A=\mathbb{P}^{n}$ and $A=\mathbb{P}^{m}\times\mathbb{P}^{n}$ [Y22]. We conjecture $X$ is algebraically hyperbolic when $\dim A\geq 5$ (see Conjecture 4.6.)
@@
    • If $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k-1)}{2}+n-3k-1$, then a very general hypersurface of $OG(k,n)$ of degree $d$ is algebraically hyperbolic.
    • If $d\leq\frac{k(2n-3k-1)}{2}+n-3k-3$, then a general hypersurface of degree $d$ of $OG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@
    • If $\dim A=\frac{k(2n-3k+1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k+1)}{2}+n+3k$, then a very general degree $d$ hypersurface of $SG(k,n)$ is algebraically hyperbolic.
    • If $d\leq\frac{k(2n-3k+1)}{2}+n+3k-2$, then a general degree $d$ hypersurface of $SG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@
    • If $\dim A=\sum_{i=0}^{m}k_{i+1}(k_{i+1}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=0}^{m}k_{j+1}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-2$ for all $1\leq i\leq m$, then a very general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
    • If $d_{i}\leq\left(\sum_{j=0}^{m}k_{j+1}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-4$ for some $i$, then a general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ contains a line. In particular, it is not algebraically hyperbolic.
Organization of the paper. In section 2, we describe the general setup we work on and obtain a first, weaker bound on algebraic hyperbolicity. Section 3 is dedicated to working out the scroll argument to the general case of homogeneous varieties and proving the bound in Theorem 1.4. In Section 4 we show when the hypersurface contains lines, thus obtaining the lower bound to algebraic hyperbolicity. In Section 5, we apply the bounds obtained in the previous sections to explicit examples of homogeneous varieties.
Acknowledgements. The author is supported by a CNPq post-doctoral fellowship. This manuscript is part of my Ph.D. thesis at the University of Illinois at Chicago. I would like to thank my advisor Izzet Coskun for our weekly meetings and support in writing this paper. I also thank Eric Riedl, Wern Yeong, and Jackson Morrow for their valuable advice and suggestions.
@@ 2.1. Setup
We first describe the general setup for the techniques developed by Clemens and Ein [Ei88, Ei91], Voisin [Vo96, Vo98], Pacienza [Pa03, Pa04], Coskun and Riedl [CR19a, CR19] and Yeong [Y22].
Let $A$ be a smooth, complex projective variety of dimension $D$, and assume $A$ admits a transitive group action by an algebraic group $G$. Let $\mathcal{E}$ a globally generated vector bundle on $A$, invariant under $G$, of rank $r<D-1$, and let $V=H^{0}(A,\mathcal{E})$. We will be particularly interested when $\mathcal{E}$ is a very ample line bundle on $A$.
@@
    $N_{h/\mathcal{X}}$ is the cokernel of the map of vertical tangent sheaves $T_{\mathcal{Y}/A}\to T_{\mathcal{X}/A}$.
@@ 2.2. The degree of the normal bundle and section-dominating line bundles
(cf [CR19, Lemma 2.2]) The following holds:
@@
By Proposition 2.1, we have a surjection $M_{\mathcal{E}}|_{Y_{t}}\to N_{h_{t}/X_{t}}$. We can get a surjection to $N_{h_{t}/X_{t}}$ from a simpler bundle by using the notion of section-dominating line bundles.
@@
(cf [CR19, Definition 2.3]) Let $\mathcal{E}$ be a vector bundle on $A$. A collection of non-trivial globally generated line bundles $L_{1},\ldots,L_{m}$ is a section-dominating collection of line bundles for $\mathcal{E}$ if $\mathcal{E}\otimes L_{i}^{\vee}$ is globally generated for every $1\leq i\leq m$ and the map
@@
Let $A=G(k,n)$ be the Grassmannians of $k$-planes in the affine $n$-space, and let $H$ correspond to the hyperplane class, that is, the pullback $p^{*}\mathcal{O}_{\mathbb{P}^{N}}(1)$ via the Plücker embedding $p:G(k,n)\to\mathbb{P}^{N}$. Then, for any $\mathcal{E}=dH$, $d>0$, $H$ is a section-dominating line bundle for $\mathcal{E}$. This follows directly from the previous example and the projective normality of the Grassmannian under the Plücker embedding (cf [RR85].) By [RR85], a similar result is true for flag varieties and Schubert varieties.
@@
(cf [CR19, Proposition 2.7]) Let $\mathcal{E}$ be a globally generated vector bundle and $M_{\mathcal{E}}$ the Lazarsfeld-Mukai bundle associated to $\mathcal{E}$. Let $L_{1},\ldots,L_{m}$ be a section-dominating collection of line bundles for $\mathcal{E}$. Then, there is a surjection
@@
From Proposition 2.7 and the surjection $M_{\mathcal{E}}|_{Y_{t}}\to N_{h_{t}/X_{t}}$, we obtain a surjection to the normal bundle
$\beta:\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\to N_{h_{t}/X_{t}}.$
By choosing generic polynomials of $H^{0}(\mathcal{E}\otimes L_{i}^{\vee})$ for the multiplication in Proposition 2.7, the rank of the image of $\beta$ is a strictly increasing function of the $s_{i}$ until surjectivity is reached (cf [Cl03].) In particular, we can choose a surjection $\beta$ so that $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{h_{t}/X_{t}}$.
@@
Let $\beta:\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\to N_{h_{t}/X_{t}}$ be a surjective map as above. Then
@@
Then by the definition of the $M_{L_{i}}$, we get an injection $K\to\bigoplus_{i=1}^{m}H^{0}(A,L_{i})\otimes\mathcal{O}_{A}|_{Y_{t}}$. Since it is a direct sum of trivial bundles, $\bigoplus_{i=1}^{m}H^{0}(A,L_{i})\otimes\mathcal{O}_{A}|_{Y_{t}}$ is a semi-stable vector bundle. It follows that
@@
Therefore, if we can control the integers $s_{i}$, we can prove the algebraic hyperbolicity of $X_{t}$. This leads us to our first criteria for hyperbolicity. This result is similar to [CR19, Cor.2.9] and [HI21, Thm 3.6].
@@
Let $A$ be a homogeneous variety of dimension $D$. Let $L_{1},\ldots,L_{m}$ be a collection of section-dominating line bundles for $\mathcal{E}$ on $A$. If, for some $\epsilon>0$,
@@
This initial setup can also be defined when $A$ is not homogeneous but contains a Zariski-open homogeneous set $A_{0}$. Results equivalent to Proposition 2.9 can be obtained in this case, see for example, [CR19, Theorem 1.2]. This allowed Coskun and Riedl [CR19] to classify algebraically hyperbolic surfaces $X$ in the cases when $A=$ a Hirzebruch surface $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(e))$, $A=\mathbb{P}^{3}$ blown up at a single point, and $A=$ a weighted projective space $\mathbb{P}(1,1,1,n)$. We believe the generalizations made in this paper can also be worked out when $A$ contains a Zariski-open homogeneous set $A_{0}$, allowing future studies of higher dimensional toric varieties, such as blowups of projective spaces and weighted projective spaces.

### v3

# Algebraic Hyperbolicity of Very General Hypersurfaces in Homogeneous Varieties

@@ Abstract.
The author is supported by CNPq Post-doctoral fellowship 152039/2024-4
@@ 1. Introduction
is satisfied, where $g(C)$ is the geometric genus of $C$. The goal of this paper is to prove the algebraic hyperbolicity of very general hypersurfaces in homogeneous varieties by generalizing techniques developed by Clemens [Cle86, Cle03], Ein [Ein88, Ein91], Voisin [Voi96, Voi98], Pacienza [Pac03, Pac04], Clemens and Ran [CR04], Coskun and Riedl [CR19, CR23] and Yeong [Yeo25].
A complex manifold $X$ is called (Brody) hyperbolic if every entire map $f:\mathbb{C}\to X$ is constant, and it is Kobayashi hyperbolic if the Kobayashi pseudometric is nondegenerate. In general, Kobayashi hyperbolicity implies Brody hyperbolicity, and when $X$ is a compact complex manifold, Brody proved that the converse holds [Bro78]. Hyperbolicity is conjectured to have deep connections to the geometry and arithmetic of varieties:
@@
See Demailly’s survey paper [Dem20] for an exposition on hyperbolicity.
It can, however, be challenging to prove the hyperbolicity of a given variety. Demailly introduced algebraic hyperbolicity as an algebraic analogue for hyperbolicity, and proved that, for smooth projective varieties, hyperbolicity implies algebraic hyperbolicity and conjectured the converse [Dem97].
A lot of progress has been made in the study of the algebraic hyperbolicity of very general hypersurfaces in $\mathbb{P}^{n}$. For $n=3$, Xu [Xu94] proved that very general hypersurfaces $X\subset\mathbb{P}^{3}$ of degree $d\geq 6$ are algebraically hyperbolic. This was improved by Coskun and Riedl [CR19], who showed that a very general quintic surface is also algebraically hyperbolic. Since surfaces of degree at most four contain rational curves and cannot be algebraically hyperbolic, the classification for very general hypersurfaces in $\mathbb{P}^{3}$ is complete. For $n\geq 4$, Clemens and Ein [Cle86, Ein88] proved that $X$ is algebraically hyperbolic when $d\geq 2n$. This was improved to $d\geq 2n-1$ by Voisin [Voi96, Voi98]. Pacienza [Pac04] and Clemens and Ran [CR04] proved it for $d\geq 2n-2$ and $n\geq 6$. Yeong [Yeo25] did the octic 4-fold case and improved the result to $d\geq 2n-2$ and $n\geq 5$. If $d\leq 2n-3$, $X$ contains lines, thus it is not algebraically hyperbolic. Hence, the only remaining case in $\mathbb{P}^{n}$ is for sextic threefolds.
Haase and Ilten [HI21] started to classify algebraically hyperbolic surfaces in toric threefolds, and gave almost sharp bounds for the cases $\mathbb{P}^{2}\times\mathbb{P}^{1},\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$, the blowup of $\mathbb{P}^{3}$ at a point, and the weighted projective space $\mathbb{P}(1,1,1,n)$. Coskun and Riedl [CR23] streamlined their scroll method and used a careful analysis of where various line bundles fail to be section-dominating to obtain sharp bounds in the cases $\mathbb{P}^{2}\times\mathbb{P}^{1},\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$, the blowup of $\mathbb{P}^{3}$ at a point, and $\mathbb{F}_{e}\times\mathbb{P}^{1}$, and partial bounds for very general surfaces in other threefolds. Robins [Rob23] uses Haase and Ilten’s method to determine many algebraically hyperbolic surfaces in toric varieties with Picard rank 2 and 3. Yeong [Yeo25] applies Coskun and Riedl’s techniques to almost completely classify algebraically hyperbolic hypersurfaces in $\mathbb{P}^{m}\times\mathbb{P}^{n}$, except for some degrees in $\mathbb{P}^{3}\times\mathbb{P}^{1}$.
@@
Let $A$ be a smooth complex projective variety such that:
  • $A$ is homogeneous with a transitive action by an algebraic group $G$;
  • $A$ has an embedding into a product of projective spaces
    $i:A\hookrightarrow\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$
    such that the divisors $H_{1},\ldots,H_{m}$ of $A$ corresponding to the pull-backs of the hyperplane classes $\mathcal{O}_{\mathbb{P}^{N_{i}}}(1)$ generate the Picard group of $A$;
  • The embedding $i$ is projectively normal, that is, the restriction
    $H^{0}\left(\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}},\mathcal{O}_{\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}}(d_{1},\ldots,d_{m})\right)\to H^{0}(A,d_{1}H_{1}+\cdots+d_{m}H_{m})$
    is surjective for all $d_{1},\ldots,d_{m}>0$.
Write the class of the canonical divisor of $A$ as $K_{A}=a_{1}H_{1}+\cdots+a_{m}H_{m}$. Let $\mathcal{E}$ be the line bundle $d_{1}H_{1}+\cdots+d_{m}H_{m}$ for $d_{1},\ldots,d_{m}>0$, equivariant under $G$, and assume that $H_{1},\ldots,H_{m}$ is a collection of section-dominating line bundles for $\mathcal{E}$ (see Definition 2.3). A degree $(d_{1},\ldots,d_{m})$ hypersurface $X$ of $A$ is a section of $\mathcal{E}$.
@@
The cases with $d_{i}\geq\dim A-a_{i}-3$ for all $1\leq i\leq m$ and $d_{i}=\dim A-a_{i}-3$ for some $1\leq i\leq m$ remain open. When $\dim A\leq 4$, there are examples when it fails to imply algebraic hyperbolicity (see Examples 4.4 and 4.5.) For $A=\mathbb{P}^{4}$, the case of sextic threefolds remains open. However, when $\dim A\geq 5$, it does imply algebraic hyperbolicity for $A=\mathbb{P}^{n}$ and $A=\mathbb{P}^{m}\times\mathbb{P}^{n}$ [Yeo25]. We conjecture $X$ is algebraically hyperbolic in these cases when $\dim A\geq 5$ (see Conjecture 4.6.)
@@
    • If $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k-1)}{2}+n-k-3$, then a very general hypersurface of $OG(k,n)$ of degree $d$ is algebraically hyperbolic.
    • If $d\leq\frac{k(2n-3k-1)}{2}+n-k-5$, then a general hypersurface of degree $d$ of $OG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@
    • If $\dim A=\frac{k(2n-3k+1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k+1)}{2}+n-k-1$, then a very general degree $d$ hypersurface of $SG(k,n)$ is algebraically hyperbolic.
    • If $d\leq\frac{k(2n-3k+1)}{2}+n-k-3$, then a general degree $d$ hypersurface of $SG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
@@
    • If $\dim A=\sum_{i=1}^{m}k_{i}(k_{i+1}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-2$ for all $1\leq i\leq m$, then a very general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
    • If $d_{i}\leq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-4$ for some $i$, then a general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ contains a line. In particular, it is not algebraically hyperbolic.
Organization of the paper
In section 2, we describe the general setup we work on and obtain a first, weaker bound on algebraic hyperbolicity. Section 3 is dedicated to working out the scroll argument to homogeneous varieties and in proving the algebraic hyperbolicity statement in Theorem 1.4. In Section 4, we determine when general hypersurfaces in homogeneous varieties contain lines, and conclude the non-algebraically hyperbolic cases from Theorem 1.4. In Section 5, we apply the bound obtained in the previous sections to specific examples of homogeneous varieties.
Acknowledgements
This manuscript is part of my Ph.D. thesis at the University of Illinois at Chicago. I would like to thank my advisor, Izzet Coskun, for our weekly meetings and support in writing this paper. I am grateful to Eric Riedl, Wern Yeong, and Jackson Morrow for their valuable advice and suggestions. I also want to thank the journal referees for identifying some mistakes and suggesting improvements.
@@ 2.1. Setup
We first describe the general setup for the techniques developed by Clemens and Ein [Ein88, Ein91], Voisin [Voi96, Voi98], Pacienza [Pac03, Pac04], Coskun and Riedl [CR19, CR23], and Yeong [Yeo25].
Let $A$ be a smooth, complex projective variety of dimension $D$, and assume $A$ admits a transitive group action by an algebraic group $G$. Let $\mathcal{E}$ a globally generated vector bundle on $A$, equivariant under $G$, of rank $r<D-1$, and let $V=H^{0}(A,\mathcal{E})$. We will be particularly interested when $\mathcal{E}$ is a very ample line bundle on $A$.
@@
    $N_{h/\mathcal{X}}$ is the cokernel of the map of vertical tangent sheaves $T_{\mathcal{Y}/A}\to h^{*}T_{\mathcal{X}/A}$.
@@ 2.2. The degree of the normal bundle and section-dominating line bundles
(cf. [CR23, Lemma 2.2]) The following holds:
@@
By Proposition 2.1, we have a surjection
(2.1) $M_{\mathcal{E}}|_{Y_{t}}\to N_{h_{t}/X_{t}}.$
We can get a surjection to $N_{h_{t}/X_{t}}$ from a simpler bundle by using the notion of section-dominating line bundles.
@@
(cf. [CR23, Definition 2.3]) Let $\mathcal{E}$ be a vector bundle on $A$. A collection of non-trivial globally generated line bundles $L_{1},\ldots,L_{m}$ is a section-dominating collection of line bundles for $\mathcal{E}$ if $\mathcal{E}\otimes L_{i}^{\vee}$ is globally generated for every $1\leq i\leq m$ and the map
@@
Let $A=G(k,n)$ be the Grassmannians of $k$-planes in the affine $n$-space, and let $H$ correspond to the hyperplane class, that is, the pullback $p^{*}\mathcal{O}_{\mathbb{P}^{N}}(1)$ via the Plücker embedding $p:G(k,n)\to\mathbb{P}^{N}$. Then, for any $\mathcal{E}=dH$, $d>0$, $H$ is a section-dominating line bundle for $\mathcal{E}$. This follows directly from the previous example and the projective normality of the Grassmannian under the Plücker embedding (cf. [RR85].) By [RR85], a similar result is true for flag varieties and Schubert varieties.
@@
(cf. [CR23, Proposition 2.7]) Let $\mathcal{E}$ be a globally generated vector bundle and $M_{\mathcal{E}}$ the Lazarsfeld-Mukai bundle associated to $\mathcal{E}$. Let $L_{1},\ldots,L_{m}$ be a section-dominating collection of line bundles for $\mathcal{E}$. Then, there is a surjection
@@
From Proposition 2.7 and the surjection (2.1), we obtain a surjection to the normal bundle
(2.2) $\beta:\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\to N_{h_{t}/X_{t}}.$
By choosing generic sections of $H^{0}(\mathcal{E}\otimes L_{i}^{\vee})$ for the multiplication in Proposition 2.7, the rank of the image of $\beta$ is a strictly increasing function of the $s_{i}$ until surjectivity is reached (cf. [Cle03, Section 2.2].) In particular, we can choose a surjection $\beta$ so that $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{h_{t}/X_{t}}$.
@@
Let $\beta:\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\to N_{h_{t}/X_{t}}$ be a surjective map as in (2.2). Then
@@
Then by the definition of the $M_{L_{i}}$, we get an injection $K\to\bigoplus_{i=1}^{m}H^{0}(A,L_{i})^{\oplus s_{i}}\otimes\mathcal{O}_{A}|_{Y_{t}}$. Since it is a direct sum of trivial bundles, $\bigoplus_{i=1}^{m}H^{0}(A,L_{i})^{\oplus s_{i}}\otimes\mathcal{O}_{A}|_{Y_{t}}$ is a semi-stable vector bundle. It follows that
@@
Therefore, if we can control the integers $s_{i}$, we can prove the algebraic hyperbolicity of $X_{t}$. This leads us to our first criterion for hyperbolicity. This result is similar to [CR23, Cor.2.9] and [HI21, Thm 3.6].
@@
Let $A$ be a homogeneous variety of dimension $D$ and let $\mathcal{E}$ be a very ample line bundle on $A$ as in Setup 2.1. Let $L_{1},\ldots,L_{m}$ be a collection of section-dominating line bundles for $\mathcal{E}$ on $A$. If, for some $\epsilon>0$,
@@
This initial setup can also be defined when $A$ is not homogeneous but contains a Zariski-open homogeneous set $A_{0}$. Results equivalent to Proposition 2.9 can be obtained in this case, see for example, [CR23, Theorem 1.2]. This allowed Coskun and Riedl [CR23] to classify algebraically hyperbolic surfaces $X$ in the cases when $A$ is a Hirzebruch surface $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(e))$, $A$ is $\mathbb{P}^{3}$ blown up at a single point, and $A$ is a weighted projective space $\mathbb{P}(1,1,1,n)$. We believe the generalizations made in this paper can also be worked out when $A$ contains a Zariski-open homogeneous set $A_{0}$, allowing future studies of higher-dimensional toric varieties, such as blowups of projective spaces and weighted projective spaces.